OpenStudy (anonymous):

If g(x) = x^2 f(x), f(4) = 3 and f'(4) = -5, then g'(4) =?

OpenStudy (anonymous):

differentaite both sides you g'(x)=2xf(x)+x^2f'(x) (by product rule) g'(4)=8f(x)+16f'(x)=24-80=-56

OpenStudy (anonymous):

You know product rule right?

OpenStudy (anonymous):

f'g + fg'

OpenStudy (anonymous):

yeah so just apply it here g=x^2 and f is f only

OpenStudy (anonymous):

wait y do u have to use g=x^2 and f =f ...didnt u answer te hquestion already and get -56

OpenStudy (anonymous):

err yeah i was just explaining how i got the answer \m/

OpenStudy (anonymous):

im still kinda confused

OpenStudy (anonymous):

after u differentiated

OpenStudy (anonymous):

If you want to get the answer just use product rule differentiate and get substitue the given values where exactly are u confused?

OpenStudy (anonymous):

oh ok nvr mind....i was confused with how u got thsoe numbers, i forgot to substite

OpenStudy (anonymous):

after i differentaited i put x=4 as you want g'(4) you know 2x=8,f(4)=3 and f'(4)=-5 i just substitued those in the equation i got to get the answer.

OpenStudy (anonymous):

awesome...thanks! =) i get the problems after i see how to approach it

OpenStudy (anonymous):

if i post anotehr problem on teh side can u help me

OpenStudy (anonymous):

Yeah sure....